using the balance sheet equation calculate

1.strand corporation recivied a copy right cost 200000 7 year legal life span 5 year estimated useful life span.

2.calculate and record amortization expense for year 2013 2014
Answer #1

Amortization expense

For 2013 = 200,000/5 years = 40,000

For 2014 = 200,000/5 years = 40,000

Dec 31,2013 Amortization expense - Copyright 40,000
Accumulated amortization - Copyright 40,000
Dec 31,2014 Amortization expense - Copyright 40,000
Accumulated amortization - Copyright 40,000